### v1

# Rearranging small sets for distinct partial sums

Abstract.
A conjecture of Graham (repeated by Erdős) asserts that for any set $A\subseteq\mathbb{F}_{p}\setminus\{0\}$, there is an ordering $a_{1},\ldots,a_{|A|}$ of the elements of $A$ such that the partial sums $a_{1},a_{1}+a_{2},\ldots,a_{1}+a_{2}+\cdots+a_{|A|}$ are all distinct. We give a very short proof of this conjecture for sets $A$ of size at most $\log p/\log\log p$.
1. A conjecture of Graham
For $A$ a finite subset of an abelian group, say that an ordering $a_{1},\ldots,a_{|A|}$ of the elements of $A$ is valid if the partial sums $a_{1},a_{1}+a_{2},\ldots,a_{1}+a_{2}+\cdots+a_{|A|}$ are all distinct. The following striking conjecture first appeared in a 1971 open problem list of Ronald Graham [10] and was later repeated in a book of Erdős and Graham [8] (see also [2, Problem #475]).
Conjecture 1.1 ([10]).
Let $p$ be a prime. Then every subset $A\subseteq\mathbb{F}_{p}\setminus\{0\}$ has a valid ordering.
Although many papers have been written about this conjecture and related problems, the state of the art is still essentially that Conjecture 1.1 holds when $|A|\leq 12$ (see [7] and the references therein) and when $p-3\leq|A|\leq p-1$ (see [12] and the references therein). Many of the arguments for small $A$ use the Polynomial Method and rely on Alon’s Combinatorial Nullstellensatz.
Alspach (as attributed in [3]) independently posed a very similar conjecture for finite cyclic groups, and versions in other groups (both abelian and nonabelian) have been studied; for further history and more extensive references, see the recent papers [4, 6]. We also mention that this line of inquiry is related to combinatorial designs and the Hall–Paige Conjecture, as described in [14].
In this short note, we make modest partial progress towards Conjecture 1.1 by showing that it holds for small sets $A$; the novelty is that our bound $\log p/\log\log p$ tends to infinity with $p$.
Theorem 1.2.
Let $p$ be a prime. Then every subset $A\subseteq\mathbb{F}_{p}\setminus\{0\}$ of size
$|A|\leq\frac{\log p}{\log\log p}$
has a valid ordering.
We prove this theorem by applying a “rectification” result of of Lev [13] (refining work of Bilu, Lev, and Ruzsa [1]) and then establishing the “integer version” of Conjecture 1.1, as follows.
Theorem 1.3.
Every finite subset $A\subseteq\mathbb{Z}\setminus\{0\}$ has a valid ordering.
2. Proofs
We begin with the integer version of Conjecture 1.1.
Proof of Theorem 1.3.
We will prove the stronger statement that there is a valid ordering of $A$ in which all of the positive elements appear before all of the negative elements. Let $P,N$ be sets of positive integers such that $A=P\cup(-N)$. It suffices to find orderings $p_{1},\ldots,p_{|P|}$ of $P$ and $n_{1},\ldots,n_{|N|}$ of $N$ such that
$p_{1}+\cdots+p_{i}\neq n_{1}+\cdots+n_{j}\quad\text{unless $(i,j)\in\{(0,0),(|P|,|N|)\}$};$
then the ordering
$p_{|P|},p_{|P|-1},\ldots,p_{1},n_{1},n_{2},\ldots,n_{|N|}$
of $A$ is valid.
We proceed by induction on $|A|$, where the base case $|A|=0$ is trivial. Now suppose that $|A|\geq 1$. Without loss of generality, we have
$\sum_{p\in P}p\geq\sum_{n\in N}n.$
The desired conclusion is obvious if $|P|<2$, so suppose that $|P|\geq 2$. Then there is some $p^{*}\in P$ such that
$\sum_{p\in P\setminus\{p^{*}\}}p\neq\sum_{n\in N}n;$
let $P^{\prime}:=P\setminus\{p^{*}\}$. The induction hypothesis provides orderings $p_{1},...,p_{|P|-1}$ of $P^{\prime}$ and $n_{1},...,n_{|N|}$ of $N$ such that
$p_{1}+\cdots+p_{i}\neq n_{1}+\cdots+n_{j}\quad\text{unless $(i,j)\in\{(0,0),(|P|-1,|N|)\}$}.$
Our choice of $p^{*}$ ensures that we also have
$p_{1}+\cdots+p_{|P|-1}\neq n_{1}+\cdots+n_{|N|}.$
Now, the orderings $p_{1},...,p_{|P|-1},p^{*}$ of $P$ and $n_{1},...,n_{|N|}$ of $N$ are as desired. ∎
We turn next to the rectification result that we will use to deduce Theorem 1.2 from Theorem 1.3. Let $A,B$ be subsets of (possibly different) abelian groups. We say that a bijection $f:A\to B$ is an $\ell$-Freiman isomorphism if
$x_{1}+\cdots+x_{\ell}=y_{1}+\cdots+y_{\ell}\Longleftrightarrow f(x_{1})+\cdots+f(x_{\ell})=f(y_{1})+\cdots+f(y_{\ell})$
for all $x_{1},\ldots,x_{\ell},y_{1},\ldots,y_{\ell}\in A$ (allowing repetitions). If $0\in A$ and $B$ has no nonzero elements of order dividing $\ell$ (as when, for instance, the group containing $B$ is torsion-free), then every $\ell$-Freiman isomorphism $f$ satisfies $f(0)=0$; in this case, we conclude that $f$ is also a $k$-Freiman isomorphism for all $k<\ell$. Bilu, Lev, and Ruzsa [1] used the Pigeonhole Principle to show that small subsets of $\mathbb{F}_{p}$ are always Freiman-isomorphic, with high order, to sets of integers. We will use the following (optimal) refinement due to Lev [13].
Theorem 2.1 ([13]).
Let $\ell\in\mathbb{N}$, let $p$ be a prime, and let $A\subseteq\mathbb{F}_{p}$. If $|A|\leq\lceil\log p/\log\ell\rceil$, then $A$ is $\ell$-Freiman-isomorphic to a set of integers.
See [11] for further discussion of rectification principles in additive combinatorics.
We are now ready to prove Theorem 1.2.
Proof of Theorem 1.2.
Let $A^{\prime}:=A\cup\{0\}$. Then $|A^{\prime}|\leq\lceil\log p/\log(|A|-1)\rceil$, and Theorem 2.1 provides an $(|A|-1)$-Freiman isomorphism $f$ from $A^{\prime}$ to some set $B$ of integers. Note that $0\notin f(A)$, and that every valid ordering of $f(A)$ pulls back to a valid ordering of $A$ (since the conditions for an ordering to be valid can be described in terms of non-equalities of sums with length at most $|A|-1$). The result now follows from Theorem 1.3. ∎
3. Remarks
  1. (1)
    By replacing Theorem 2.1 with Lev’s more general rectification criterion for abelian groups [13], we can extend Theorem 1.2 to abelian groups with no elements of small torsion: If $G$ is an abelian group with no nonzero elements of order strictly smaller than $p$, then every subset $A\subseteq G\setminus\{0\}$ of size at most $\log p/\log\log p$ has a valid ordering. (See the discussion in [6, 7] for previous results in this direction.)
  2. (2)
    It is known (see, e.g., [11], following [9]) that even moderate-sized subsets of $\mathbb{F}_{p}$ can be rectified if one adds a small-doubling assumption; again, our arguments apply to this scenario.
  3. (3)
    Say that an abelian group $G$ is sequenceable if every subset of $G\setminus\{0\}$ has a valid ordering. The argument of Theorem 1.3 can be easily modified to show the stronger statement that an abelian group $H$ is sequenceable if and only if $H\times\mathbb{Z}$ is sequenceable: In the proof, consider valid orderings of $A$ in which all of the elements with positive second coordinate appear before all of the elements with second coordinate 00, which in turn appear before all of the elements with negative second coordinate. (Theorem 1.3 corresponds to the case where $H$ is the trivial group.) It could be interesting to formulate versions of this principle in nonabelian settings (see [5]).
  4. (4)
    Combining the first and third remarks, we see that if $H_{1}$ is a sequenceable abelian group and $H_{2}$ is an abelian group with no nonzero elements of order strictly smaller than $p$, then every subset $A\subseteq(H_{1}\times H_{2})\setminus\{(0,0)\}$ with $|\pi_{H_{2}}(A)|\leq\log p/\log\log p$ has a valid ordering. In the context of finite cyclic groups, this significantly extends the results of [6] and some of the results of [5].
  5. (5)
    The valid orderings constructed in our proofs of Theorems 1.3 and 1.2 are “two-sided” in the sense that their reverses are also valid; this condition is equivalent to the absence of zero-sum proper consecutive suborderings. In the setting of Conjecture 1.1 (and its generalization to other abelian groups), is it always possible to find two-sided valid orderings?
Acknowledgements
The author was supported in part by the NSF Graduate Research Fellowship Program under grant DGE–203965. I thank Stefan Steinerberger for bringing Graham’s conjecture to my attention and for engaging in helpful discussions, and I thank Simone Costa for giving useful comments on a draft of this paper.